# Calibration Method of Spacecraft–Inertial Sensor Center-of-Mass Offset for the Taiji Gravitational Wave Detection Mission under Science Mode

### v1

@@ IV.3 Simulation Result
Using the science mode readouts, the ES-AKF estimates the CoM offsets $\mathbf{b}_{1}^{O_{1}}$ and $\mathbf{b}_{2}^{O_{2}}$ of TM1 and TM2 relative to the spacecraft. The convergence behavior of these estimates is presented in Figures 11 and 12.
@@
Figures 11 and 12 show that the estimates of $\mathbf{b}_{i,x}^{O_{i}}$ exhibit slower convergence compared to those of $\mathbf{b}_{i,y}^{O_{i}}$ and $\mathbf{b}_{i,z}^{O_{i}}$. This behavior arises because the $x$-axis CoM calibration relies solely on IS readout signals ($\tilde{y}_{m_{i}O_{i}}^{O_{i}}$, $\tilde{z}_{m_{i}O_{i}}^{O_{i}}$), which have lower precision, whereas the $y$- and $z$-axis calibrations additionally utilize the high-precision IFO readout ($\tilde{x}_{m_{i}O_{i}}^{O_{i}}$). Consequently, $\mathbf{b}_{i,y}^{O_{i}}$ and $\mathbf{b}_{i,z}^{O_{i}}$ converge using data of approximately $0.4\times 10^{6}~\mathrm{s}$ in duration, while the $\mathbf{b}_{i,x}^{O_{i}}$ requires approximately $1.4$ to $1.6\times 10^{6}~\mathrm{s}$. Correspondingly, when evaluated using the confidence level determined from the error covariance matrix, the final calibration accuracy of $\mathbf{b}_{1,x}^{O_{1}}$ is lower than that of $\mathbf{b}_{1,y}^{O_{1}}$ and $\mathbf{b}_{1,z}^{O_{1}}$. Table 2 lists the calibrated values, the true values, and the confidence levels computed from the final state error covariance matrix $\mathbf{P}_{\mathrm{end}}$ (taken as $\mathbf{P}_{\infty}$) at the termination time.
@@
$\mathbf{b}_{1,x}^{0}$ $9.081\text{\times}{10}^{-2}$ $1.335\text{\times}{10}^{-3}$ $9.087\text{\times}{10}^{-2}$
$\mathbf{b}_{1,y}^{0}$ $1.540\text{\times}{10}^{-1}$ $1.705\text{\times}{10}^{-4}$ $1.554\text{\times}{10}^{-1}$
$\mathbf{b}_{1,z}^{0}$ $5.086\text{\times}{10}^{-4}$ $9.219\text{\times}{10}^{-5}$ 0
$\mathbf{b}_{2,x}^{0}$ $9.019\text{\times}{10}^{-2}$ $3.762\text{\times}{10}^{-3}$ $9.087\text{\times}{10}^{-2}$
$\mathbf{b}_{2,y}^{0}$ $1.553\text{\times}{10}^{-1}$ $1.394\text{\times}{10}^{-4}$ $1.554\text{\times}{10}^{-1}$
$\mathbf{b}_{2,z}^{0}$ $3.509\text{\times}{10}^{-4}$ $7.746\text{\times}{10}^{-5}$ 0
Based on the results in Table 2, the system identification algorithm using the ES-AKF estimator achieves high-precision calibration of the SC-TM CoM offset in science mode using only IS/IFO/DWS readout signals. The confidence levels are better than $4~\mathrm{mm}$ for $\mathbf{b}_{i,x}^{O_{i}}$, better than $0.2~\mathrm{mm}$ for $\mathbf{b}_{i,y}^{O_{i}}$, and better than $0.1~\mathrm{mm}$ for $\mathbf{b}_{i,z}^{O_{i}}$. However, the calibrated value of $\mathbf{b}_{i,x}^{O_{i}}$ is actually closer to the true value than that of $\mathbf{b}_{i,y}^{O_{i}}$. This phenomenon can be explained by two factors.
First, the calibration value is determined as the mathematical expectation of all estimates over the final $3600~\mathrm{s}$ of data. In theory, the estimate of $\mathbf{b}_{i}^{O_{i}}$ oscillates around or slowly approaches the true value. Therefore, the close agreement between the $\mathbf{b}_{i,x}^{O_{i}}$ estimate and the true value indicates that its convergence point is near the true value. Nevertheless, its error variance—i.e., the amplitude of oscillation—remains the largest among the three degrees of freedom. This behavior is clearly illustrated by the convergence curves in Figures 11 and 12.
Second, the relative displacement $\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$ exhibits higher noise in the ultra-low-frequency band below the control bandwidth along the sensitive axis (as shown in Figure 5), which does not affect control requirements but significantly impacts the estimation accuracy of $\mathbf{x}_{m_{i}O_{i}}^{O_{i}}$ by the AKF in the ultra-low-frequency range (see Figure 10), thereby affecting the estimation of $\mathbf{b}_{i,y}^{O_{i}}$.
@@ V Conclusion
This work proposes and validates a maneuver-free calibration scheme for the SC-TM CoM offset during the science operation of space-based gravitational-wave antennas, especially for Taiji (a typical LISA-like mission). The method directly utilizes the standard science-mode readouts of the IS, interferometer, and differential wavefront sensor, eliminating the need for dedicated spacecraft maneuvers that disrupt measurement continuity and laser-link stability.
A high-fidelity, closed-loop simulation framework was developed, integrating a detailed nonlinear dynamics model of the LISA-like SC-TM system and a robust $H_{\infty}$ drag-free and attitude controller. Using 25 days of simulated science data, we designed an Extended-State Adaptive Kalman Filter that incorporates the CoM offset induced acceleration noises, sensor noise dynamics, and systematic biases into a unified state-estimation framework. The filter successfully identifies the CoM offset vector with an accuracy of 0.01–1.5 mm across all axes, corresponding to a maximum relative error below 1%.
Beyond calibration, the proposed ES-AKF demonstrates the capability of high-precision filtering of the IS readout signal and suppression of noise. Under the condition that the IS noise is accurately modeled, the filter suppresses the IS readout noise from $1\times 10^{-9}~\mathrm{m}/\sqrt{\mathrm{Hz}}$ to $1\times 10^{-11}~\mathrm{m}/\sqrt{\mathrm{Hz}}$. This result suggests that the ES-AKF can be integrated or embedded into the DFACS to enhance measurement quality and significantly improve control performance.

### v2

@@ IV.3 Simulation Result
Using the science mode readouts, the ES-AKF estimates the CoM offsets $\mathbf{b}_{1}^{O_{1}}$ and $\mathbf{b}_{2}^{O_{2}}$ of TM1 and TM2 relative to the SC. The convergence behavior of these estimates is presented in Figs. 10 and 11.
@@
Figs. 10 and 11 show that the estimates of $\mathbf{b}_{i,x}^{O_{i}}$ exhibit slower convergence compared to those of $\mathbf{b}_{i,y}^{O_{i}}$ and $\mathbf{b}_{i,z}^{O_{i}}$. This behavior arises because the $x$-axis CoM calibration relies solely on IS readout signals ($\tilde{y}_{m_{i}O_{i}}^{O_{i}}$, $\tilde{z}_{m_{i}O_{i}}^{O_{i}}$), which have lower precision, whereas the $y$- and $z$-axis calibrations additionally utilize the high-precision IFO readout ($\tilde{x}_{m_{i}O_{i}}^{O_{i}}$). Consequently, $\mathbf{b}_{i,y}^{O_{i}}$ and $\mathbf{b}_{i,z}^{O_{i}}$ converge using data of approximately $0.4\times 10^{6}~\mathrm{s}$ in duration, while the $\mathbf{b}_{i,x}^{O_{i}}$ requires approximately $1.4$ to $1.6\times 10^{6}~\mathrm{s}$. Correspondingly, for the confidence level determined from the error covariance matrix, the final calibration accuracy of $\mathbf{b}_{1,x}^{O_{1}}$ is lower than that of $\mathbf{b}_{1,y}^{O_{1}}$ and $\mathbf{b}_{1,z}^{O_{1}}$. Table 3 lists the calibrated values, the true values, and the confidence levels computed from the final state error covariance matrix $\mathbf{P}_{\mathrm{end}}$ (taken as $\mathbf{P}_{\infty}$) at the termination time.
@@
$\mathbf{b}_{1,x}^{O_{1}}$ $9.081\text{\times}{10}^{-2}$ $1.335\text{\times}{10}^{-3}$ $9.087\text{\times}{10}^{-2}$
$\mathbf{b}_{1,y}^{O_{1}}$ $1.540\text{\times}{10}^{-1}$ $1.705\text{\times}{10}^{-4}$ $1.554\text{\times}{10}^{-1}$
$\mathbf{b}_{1,z}^{O_{1}}$ $5.086\text{\times}{10}^{-4}$ $9.219\text{\times}{10}^{-5}$ 0
$\mathbf{b}_{2,x}^{O_{2}}$ $9.019\text{\times}{10}^{-2}$ $3.762\text{\times}{10}^{-3}$ $9.087\text{\times}{10}^{-2}$
$\mathbf{b}_{2,y}^{O_{2}}$ $1.553\text{\times}{10}^{-1}$ $1.394\text{\times}{10}^{-4}$ $1.554\text{\times}{10}^{-1}$
$\mathbf{b}_{2,z}^{O_{2}}$ $3.509\text{\times}{10}^{-4}$ $7.746\text{\times}{10}^{-5}$ 0
Based on the results in Table 3, the system identification algorithm using the ES-AKF estimator achieves high-precision calibration of the SC-TM CoM offset in science mode using only IS/IFO/DWS readout signals. The confidence levels are better than $4~\mathrm{mm}$ for $\mathbf{b}_{i,x}^{O_{i}}$, better than $0.2~\mathrm{mm}$ for $\mathbf{b}_{i,y}^{O_{i}}$, and better than $0.1~\mathrm{mm}$ for $\mathbf{b}_{i,z}^{O_{i}}$. However, the calibrated value of $\mathbf{b}_{i,x}^{O_{i}}$ is actually closer to the true value than that of $\mathbf{b}_{i,y}^{O_{i}}$. This can be explained by the following two reasons.
First, the calibration value is determined as the mathematical expectation of all estimates over the final $3600~\mathrm{s}$ of data. In theory, the estimate of $\mathbf{b}_{i}^{O_{i}}$ oscillates around or slowly approaches the true value. Therefore, the close agreement between the $\mathbf{b}_{i,x}^{O_{i}}$ estimate and the true value indicates that its convergence point is near the true value. Nevertheless, its error variance—i.e., the amplitude of oscillation—remains the largest among the three degrees of freedom. This behavior is clearly illustrated by the convergence curves in Figs. 10 and 11.
Second, the relative displacement $\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$ exhibits higher noise in the ultra-low-frequency band below the control bandwidth along the sensitive axis (as shown in Fig. 5), which does not affect control requirements but significantly impacts the estimation accuracy of $\mathbf{x}_{m_{i}O_{i}}^{O_{i}}$ by the AKF in the ultra-low-frequency range (see Fig. 9), thereby affecting the estimation of $\mathbf{b}_{i,y}^{O_{i}}$.
@@ V Conclusion
This work proposes and validates a maneuver-free calibration scheme for the SC-TM CoM offset during the science operation of space-based GW antennas, especially for Taiji (a typical LISA-like mission). The method directly utilizes the standard science-mode readouts of the IS, interferometer, and differential wavefront sensor, eliminating the need for dedicated spacecraft maneuvers that disrupt measurement continuity and laser-link stability.
A high-fidelity, closed-loop simulation framework was developed, integrating a detailed nonlinear dynamical model of the LISA-like SC-TM system and a robust $H_{\infty}$ drag-free and attitude controller. Using 25 days of simulated science data, we designed an Extended-State Adaptive Kalman Filter that incorporates the CoM offset induced acceleration noises, sensor noise dynamics, and systematic biases into a unified state-estimation framework. The filter successfully identifies the CoM offset vector with an accuracy of 0.01–1.5 mm across all axes, corresponding to a maximum relative error below 1%. Beyond calibration, the proposed ES-AKF demonstrates the capability of high-precision filtering of the IS readout signal and suppression of noise.
@@
Acknowledgements
This work was supportedin partby the National Key R&D Program of China under Grant 2021YFC2202900, in part by the National Natural Science Foundation of China under Grant 62203145 and U25B2052, in part by the Heilongjiang Provincial Natural Science Foundation under Grant YQ2024F010, and in part by the International Partnership Program of the Chinese Academy of Sciences, Grant No. 025GJHZ2023106GC.